OpenStudy (anonymous):

Is the given number a solution of the equation? 13 – d = 8; 5

OpenStudy (anonymous):

try to solve the eqn ... u can find the answer

OpenStudy (turingtest):

so you already know the answer apparently

OpenStudy (turingtest):

then the problem is 13-d=8 and you want to know the steps?

OpenStudy (merchandize):

13-d=8 -d=8-13 minus will get cancelled fom both the sides d=5 yea its correct

OpenStudy (turingtest):

or equally valid approach: 13-d=8 add d to both sides: 13=8+d subtract 8 from both sides 5=d and low and behold you get the same answer